# A Novel Terabit Grid-of-Beam Optical Wireless Multi-User Access Network With Beam Clustering

### v1

@@ III-D Design Examples
[…]
@@ IV-A Coordinated Multi-Beam Transmission
For beam clustering based on CoMP, three main categories are identified [23]: static clustering, semi-dynamic clustering, and dynamic clustering. In static clustering, once clusters are formed, they do not change with changes in the network. Static clusters are designed based on the network topology such as the geometry of the environment, the number of APs and their positions [23]. Semi-dynamic clustering is an enhanced version of static clustering where several layers of static clusters are formed to avoid inter-cluster interference (ICI) and users are associated with the best serving cluster. While semi-static clustering can improve potential performance gains compared to static clustering, the overlapping nature of clusters increases the system complexity [23]. In dynamic clustering, the system continuously adapts its clustering strategy to any changes in the network such as user mobility, inactive beams or load variations. Although with this scheme ICI can be minimized by dynamically adjusting cluster sizes for individual users, it comes with a higher complexity due to optimal scheduling.
The implementation complexity of CoMB increases with the number of coordinated beams and it can only be realized for a small cluster of beams in practice. The key for maximizing the benefits of CoMB lies in the optimum formation of clusters. By comparison, static clustering incurs significantly less complexity and is a suitable candidate for initial CoMB deployment in conjunction with the proposed double-tier AP.
@@ IV-B Downlink System Modeling
Fig. 1 depicts the downlink system configuration where only the middle transmitter element of the double-tier AP is shown for convenience. There are a total of $K$ user devices distributed uniformly over a horizontal communication plane, and $h_{\mathrm{DL}}$ represents the vertical separation between the AP and user devices. The coverage area of the network is partitioned into $N_{\mathrm{c}}$ non-overlapping regions, and each region is covered by an exclusive cluster of beams. According to CoMP-JT, each user device is assumed to be served by one cluster within a predefined region that provides the highest SINR. The coordinated VCSELs jointly send multiple copies of the same signal for each user. By using intensity modulation and direct detection (IM/DD), these signals are then detected at the same time by the receiver, and therefore their received powers are combined. Each user device is equipped with an ADR, the details of which are provided as follows.
@@ IV-B1 Angle Diversity Receiver
A multi-element non-imaging ADR based on compound parabolic concentrators is adopted from [17], where we put forward the optimized multi-tier ADR design under the constraints on the overall field of view (FOV) and dimensions. The rationale behind choosing this receiver design is that the use of CPCs allows the maximum concentration gain to be achieved due to the law of conservation of Etendue [24], and it offers a robust performance by relieving strict alignment requirements and enabling connectivity to multiple light beams [25, 26]. We consider a single-tier ADR composed of seven identical elements oriented in the desired spatial directions, as shown in Fig. 8. Each receiver element comprises a dielectric CPC coupled with a 2D array of PIN PDs. A two-stage array processing is performed by applying equal gain combining (EGC) to the output signals of each PD array and maximum ratio combining (MRC) to the output signals of the receiver elements. The ADR parameters include $\theta_{\mathrm{CPC}}$, $A_{\mathrm{PD}}$ and $N_{\mathrm{PD}}$, which denote the CPC acceptance angle, the PD area, and the number of PDs on the array, respectively. Also, $A_{1}$ and $A_{2}$ represent the areas of the CPC entrance and exit apertures, respectively. For each receiver element [17]:
@@
Figure 8: Downlink system model representing the optical channel gains $H_{kj}^{(i)}$ from the $i$th VCSEL to the $j$th element of the ADR, $j=1,2,\dots,7$, for the $k$th user. Note that only the central element of the double-tier AP is shown.
@@ IV-B2 Optical Channel Gain
We use the coordinate system shown in Fig. 8 for analysis. The $k$th user device is located at $\mathbf{p}_{k}=[x_{k}~{}y_{k}~{}0]^{T}$. The distance vector from the $v$th transmitter element of the AP to $\mathbf{p}_{k}$ is given by $\mathbf{d}_{kv}=\mathbf{p}_{k}-\mathbf{q}_{v}$, and:
@@
The incident beam intensity of the $i$th VCSEL with a radiance angle of $\phi_{ki}$ with respect to the user position is:
@@
where $\phi_{ki}=\cos^{-1}\left(\dfrac{\mathbf{d}_{kv}{\cdot}\mathbf{v}^{\prime}_{i}}{d_{kv}}\right)$, for $v=\Big{\lceil}\dfrac{i}{25}\Big{\rceil}$. Therefore, the optical power received by the $j$th ADR element of the user device can be expressed as:
@@
where $\psi_{ji}=\cos^{-1}(\mathbf{e}_{j}{\cdot}\mathbf{v}^{\prime}_{i})$ is the incidence angle of the beam with respect to the normal vector of the $j$th receiver element. The last factor in (40) is an indicator function which is defined as $\mathds{1}_{\theta_{\mathrm{CPC}}}(\psi_{ji})=1$ if $0\leq\psi_{ji}\leq\theta_{\mathrm{CPC}}$, and 00 otherwise. By using (35)–(40), the direct current (DC) gain of the downlink channel between the $i$th VCSEL and the $j$th receiver element of the $k$th user device can be derived as:
@@
Let $\mathcal{C}_{u}$ represent the index set of the VCSELs participating in the $u$th cluster for $u=1,2,\dots,N_{\mathrm{c}}$. Once the users joined their clusters, the $u$th cluster serves a subset of the users with $K_{u}$ members such that $\sum\limits_{u=1}^{N_{\mathrm{c}}}K_{u}=K$. The users in the network are globally indexed from $1$ to $K$ and those served by the $u$th cluster are indicated by the index set $\mathcal{K}_{u}$ with $\lvert\mathcal{K}_{u}\rvert=K_{u}$. The optical power received by the $j$th receiver element of the $k$th user from the $u$th cluster and the corresponding total received power are given by:
@@
where $H_{k}^{\mathcal{C}_{u}}$ is the equivalent DC gain of the CoMB-JT channel between the $u$th cluster and the $k$th user such that:
@@ IV-B3 Noise Model
Noise sources at the receiver consist of the thermal noise and shot noise, and the laser noise arising from the VCSELs. For the $j$th receiver element of the $k$th user, the total noise is modeled as a white Gaussian noise with a single-sided power spectral density (PSD) of [13]:
@@ IV-C Multi-User Performance Analysis
To ensure a high spectral efficiency, we assume the use of DCO-OFDM based on an $N$-point fast Fourier transform (FFT) with adaptive quadrature amplitude modulation (QAM). Thus, $\frac{N}{2}-1$ data-carrying subcarriers are available for each VCSEL. The average electrical power, $P_{\mathrm{elec}}$, is related to the average transmit optical power, $P_{\mathrm{t}}$, by $P_{\mathrm{elec}}=\frac{1}{9}P_{\mathrm{t}}^{2}$, so that the clipping distortion remains negligible [29]. We apply NOMA and OFDMA techniques for the downlink multi-user access. The received SINR and achievable rate expressions for these techniques are provided in the following.
@@ IV-C1 Multi-User Access
In NOMA, spectral resources (i.e., orthogonal frequency division multiplexing (OFDM) subcarriers) are fully shared by every user in a cluster and downlink transmissions for different users are multiplexed in the power domain based on superposition coding. In principle, for each cluster, the superposed signal is broadcast to all the involved users and successive interference cancellation (SIC) is applied at the receiver side for multi-user detection. For each user, SIC is performed by detecting and canceling the interference caused by other users with weaker channel conditions while treating the interference originating from those with stronger channel conditions as noise.
@@
Without loss of generality, suppose that the channel gains of the users associated with the $u$th cluster are sorted in an ascending order such that:
@@
for $u=1,2,\dots,N_{\mathrm{c}}$. Let $a_{k}^{\mathcal{C}_{u}}$ be the optical power allocation coefficient corresponding to the $k$th user in the $u$th cluster. The total power constraint implies that $\sum\limits_{k=1}^{K_{u}}\left(a_{k}^{\mathcal{C}_{u}}\right)^{\!2}=1$. To maintain fairness among users, a suitable power allocation policy is one that prioritizes the users that have weaker channel conditions. In this case, it follows from (45) that $a_{1}^{\mathcal{C}_{u}}\geq\cdots\geq a_{k}^{\mathcal{C}_{u}}\geq\cdots\geq a_{K_{u}}^{\mathcal{C}_{u}}$. The power allocation coefficients are chosen as $a_{k}^{\mathcal{C}_{u}}=\sqrt{\dfrac{K_{u}-k+1}{\sigma}}$, where $\sigma=\dfrac{K_{u}(K_{u}+1)}{2}$ [30]. Upon removing the DC term, the received signal at the $j$th receiver element of the $k$th user in the $u$th cluster is expressed as (46), given at the top of the page, where the message symbols $X_{\ell}^{\mathcal{C}_{v}}$ are assumed to be drawn from a normalized QAM constellation, and the factor $\zeta=\sqrt{\dfrac{N}{N-2}}$ ensures that the average power of the time domain signal is normalized to unity. Also, the noise terms $V_{kj}$ are independent Gaussian random variables with a zero mean and variance $\sigma_{kj}^{2}$, and the factor $\xi=\dfrac{N-2}{N}$ is the subcarrier utilization ratio. The received signals, for $j=1,2,\dots,7$, are combined together in the form of $Y_{k}^{\mathcal{C}_{u}}=\sum\limits_{j=1}^{7}w_{kj}Y_{kj}^{\mathcal{C}_{u}}$, where $w_{kj}$ is the corresponding combining weight. The receiver fully decodes the SIC term, and then subtracts its remodulated version from the received signal before proceeding to decode the $k$th user’s message. During the whole decoding process, the receiver treats both the residual multi-user interference (MUI) in the same cluster and the ICI from the rest of the clusters as noise. Assuming perfectly error-free SIC operations, the SINR of the $k$th user in the $u$th cluster is given by (47) at the top of next page, where $\gamma=\dfrac{\xi^{2}}{R_{\mathrm{PD}}^{2}P_{\mathrm{elec}}}$. Based on MRC, the combining weight for each receiver element is proportional to its received photocurrent to noise ratio [31], resulting in (48) at the top of the page.
When using OFDMA, a fraction $b_{k}^{\mathcal{C}_{u}}$ of the total bandwidth and a fraction $p_{k}^{\mathcal{C}_{u}}$ of the total power are allocated to the $k$th user in the $u$th cluster subject to $\sum\limits_{k=1}^{K_{u}}b_{k}^{\mathcal{C}_{u}}=1$ and $\sum\limits_{k=1}^{K_{u}}p_{k}^{\mathcal{C}_{u}}=1$. Hence, under OFDMA, the intra-cluster MUI is eliminated and the ICI remains as the residual interference. The SINR of the $k$th user in the $u$th cluster can be written as:
@@ IV-C2 Sum Rate Analysis
and $\Gamma=-\dfrac{\ln\left(5\ \!\mathrm{BER}\right)}{1.5}$ is the signal-to-noise ratio (SNR) gap accounting for the required BER performance.
The achievable rate of the $k$th user in the $u$th cluster under NOMA and OFDMA is deduced from (51) as:

### v2

@@ III-D Design Examples
We provide numerical examples to demonstrate the AP design. The results are presented by assuming a wavelength [^2] Most commercially available VCSEL devices operate in an IR wavelength range of $800-1000$ nm and they are generally cheaper than those operating at higher IR wavelengths such as $1550$ nm. Nevertheless, the choice of higher wavelengths is beneficial in terms of eye safety considerations. of $\lambda=950$ nm and a beam waist radius of $w_{0}=5$ µm. For each VCSEL array, the total size is $1$ cm2 with a pitch distance of $\delta=2$ mm, and the array-to-lens distance is $d_{\mathrm{VL}}=5$ mm. To understand the effect of the design parameters on the spatial separation between the beam spots, Figs. 6 and 6 illustrate the normalized intensity distribution of the middle transmitter element on the receiver plane at $h_{\mathrm{DL}}=3$ m underneath the AP. In Fig. 6, a relatively large lens of diameter $L=24$ mm is used. In this case, for $R_{\mathrm{lens}}=25$ mm, the beams constitute a quasi-square coverage area, and for $R_{\mathrm{lens}}=17$ mm, they are still largely overlapping. When using $R_{\mathrm{lens}}=15$ mm, despite a slight overlap between the adjacent beam spots, the beams are sufficiently separated. In Fig. 6, a smaller lens with $L=16$ mm is used. It can be observed that by reducing the radius of curvature of the lens, refraction angles increase and the beams move considerably further apart.
Aiming to achieve a full beam coverage while maintaining minimum overlaps between the beam spots, we opt for $L=16$ mm and $R_{\mathrm{lens}}=15$ mm. With this configuration, the beam spots cover the receiver plane in a complementary manner while the area covered by each of them is almost a circle of radius $w^{\prime}(z)$ (i.e., the transformed beam spot radius containing $86\%$ of the transmit power carried by the beam), as shown in Fig. LABEL:sub@Fig:LensParameters2_a. We use this configuration for a room of size $5\times 5\times 3$ m3. To extend the AP coverage over the entire room area of $25$ m2, the transmitter elements are arranged next to each other at a distance of $d_{\mathrm{lens}}=2$ cm, as shown in Fig. LABEL:sub@Fig:AP_a. To obtain the required tilt angle for each transmitter element, the beam spot center of the central VCSEL in each array is assumed to be nearly landing on the intersection of the diagonals of the corresponding quadrilateral formed at the corners of the room, and this leads to $\theta_{\mathrm{tilt}}=21^{\circ}$. Fig. 7 shows the resulting spatial distribution of the normalized intensity on the receiver plane. To examine the accuracy of the proposed analytical modeling, the results are compared with Zemax OpticStudio simulations based on non-sequential ray tracing [25]. For each VCSEL, $10^{5}$ rays are randomly generated and each ray is independently traced until it reaches the receiver plane. Comparing Figs. LABEL:sub@Fig:AP_Coverage_Analytical and LABEL:sub@Fig:AP_Coverage_Zemax, it can be seen that the analytical results and those obtained by using Zemax OpticStudio are closely matching.
@@ IV Beam Clustering for Interference Management
IV-A Coordinated Multi-Beam Transmission
For beam clustering based on CoMP, three main categories are identified [27]: static clustering, semi-dynamic clustering, and dynamic clustering. Static clustering provides a means to cancel intra-cluster interference by using a predefined set of clusters which once formed do not change with any variations in the network. Static clusters are designed based on the network topology such as the geometry of the environment, the number of APs and their positions [27]. Semi-dynamic clustering is an enhanced version of static clustering where several layers of static clusters are formed to avoid inter-cluster interference (ICI) and users are associated with the best serving cluster. While semi-static clustering can improve potential performance gains compared to static clustering, the overlapping nature of clusters increases the system complexity [27]. In dynamic clustering, the system continuously adapts its clustering strategy to any changes in the network such as user mobility, inactive beams or load variations. Although with this scheme ICI can be minimized by dynamically adjusting cluster sizes for individual users, it comes with a higher complexity due to optimal scheduling.
The implementation complexity of CoMB increases with the number of coordinated beams and it can only be realized for a small cluster of beams in practice. The key for maximizing the benefits of CoMB lies in the optimum formation of clusters. By comparison, static clustering incurs significantly less complexity and is a suitable candidate for initial CoMB deployment in conjunction with the proposed AP design. In fact, the formation of predefined clusters of beams in static clustering is compatible with the centralized architecture of the double-tier AP and it greatly facilitates the implementation of CoMB-JT as all the VCSELs are reachable at the same time.
@@ IV-B Downlink System Modeling
Fig. 8 depicts the downlink system configuration where only the middle transmitter element of the double-tier AP is shown for convenience. There are a total of $K$ UEs distributed uniformly over a horizontal communication plane. The coverage area of the network is partitioned into $N_{\mathrm{c}}$ non-overlapping regions, and each region is covered by an exclusive cluster of beams. According to CoMB-JT, each UE is assumed to be served by one cluster within a predefined region that provides the highest SINR. Coordinated VCSELs jointly send multiple copies of the same signal for each UE. By using intensity modulation and direct detection (IM/DD), these signals are then detected at the same time by the receiver, and therefore their received powers are combined. Each UE is equipped with an ADR, the details of which are provided as follows.
Figure 8: Downlink system model representing the optical channel gains $H_{kj}^{(i)}$ from the $i$th VCSEL to the $j$th element of the ADR, $j=1,2,\dots,7$, for the $k$th UE. Note that only the central element of the AP is shown.
@@ IV-B1 Angle Diversity Receiver
A multi-element non-imaging ADR based on CPCs is adopted from [17], where we put forward the optimized multi-tier ADR design under the constraints on the overall field of view (FOV) and dimensions. The rationale behind choosing this receiver design is that the use of CPCs allows the maximum concentration gain to be achieved due to the law of conservation of Etendue [28], and it offers a robust performance by relieving strict alignment requirements and enabling connectivity to multiple light beams [29, 30]. We consider a single-tier ADR composed of seven identical elements oriented in the desired spatial directions, as shown in Fig. 8. Each receiver element comprises a dielectric CPC coupled with a 2D array of PIN PDs. A two-stage array processing is performed by applying equal gain combining (EGC) to the output signals of each PD array and MRC to the output signals of the receiver elements. The ADR parameters include $\theta_{\mathrm{CPC}}$, $A_{\mathrm{PD}}$ and $N_{\mathrm{PD}}$, which denote the CPC acceptance angle, the PD area, and the number of PDs on the array, respectively. Also, $A_{1}$ and $A_{2}$ represent the areas of the CPC entrance and exit apertures, respectively. For each receiver element [17]:
@@ IV-B2 Optical Channel Gain
We use the coordinate system shown in Fig. 8 for analysis. The $k$th UE is located at $\mathbf{p}_{k}=[x_{k}~{}y_{k}~{}0]^{T}$. The distance vector from the $v$th transmitter element of the AP to $\mathbf{p}_{k}$ is given by $\mathbf{d}_{kv}=\mathbf{p}_{k}-\mathbf{q}_{v}$, and:
@@
The incident beam intensity of the $i$th VCSEL with a radiance angle of $\phi_{ki}$ with respect to the UE position is:
@@
where $\phi_{ki}=\cos^{-1}\left(\dfrac{\mathbf{d}_{kv}{\cdot}\mathbf{v}^{\prime}_{i}}{d_{kv}}\right)$, for $v=\Big{\lceil}\dfrac{i}{25}\Big{\rceil}$. Therefore, the optical power received by the $j$th ADR element of the $k$th UE can be expressed as:
@@
where $\psi_{ji}=\cos^{-1}(\mathbf{e}_{j}{\cdot}\mathbf{v}^{\prime}_{i})$ is the incidence angle of the beam with respect to the normal vector of the $j$th receiver element. The last factor in (41) is an indicator function which is defined as $\mathds{1}_{\theta_{\mathrm{CPC}}}(\psi_{ji})=1$ if $0\leq\psi_{ji}\leq\theta_{\mathrm{CPC}}$, and 00 otherwise. By using (36)–(41), the direct current (DC) gain of the downlink channel between the $i$th VCSEL and the $j$th receiver element of the $k$th UE can be derived as:
@@
Let $\mathcal{C}_{u}$ represent the index set of the VCSELs participating in the $u$th cluster for $u=1,2,\dots,N_{\mathrm{c}}$. Once the UEs joined their clusters, the $u$th cluster serves a subset of them with $K_{u}$ members such that $\sum\limits_{u=1}^{N_{\mathrm{c}}}K_{u}=K$. The UEs in the network are globally indexed from $1$ to $K$ and those served by the $u$th cluster are indicated by the index set $\mathcal{K}_{u}$ with $\lvert\mathcal{K}_{u}\rvert=K_{u}$. The optical power received by the $j$th receiver element of the $k$th UE from the $u$th cluster and the corresponding total received power are given by:
@@
where $H_{k}^{\mathcal{C}_{u}}$ is the equivalent DC gain of the CoMB-JT channel between the $u$th cluster and the $k$th UE such that:
@@ IV-B3 Noise Model
Noise sources at the receiver consist of the thermal noise and shot noise, and the laser noise arising from the VCSELs. For the $j$th receiver element of the $k$th UE, the total noise is modeled as a white Gaussian noise with a single-sided power spectral density (PSD) of [13]:
@@ IV-C Multi-User Performance Analysis
To ensure a high spectral efficiency, we assume the use of DC-biased optical orthogonal frequency division multiplexing (DCO-OFDM) based on an $N$-point fast Fourier transform (FFT) with adaptive quadrature amplitude modulation (QAM). Thus, $\frac{N}{2}-1$ data-carrying subcarriers are available for each VCSEL. The average electrical power, $P_{\mathrm{elec}}$, is related to the average transmit optical power, $P_{\mathrm{t}}$, by $P_{\mathrm{elec}}=\frac{1}{9}P_{\mathrm{t}}^{2}$, so that the clipping distortion remains negligible [33]. We apply NOMA and OFDMA techniques for the downlink multi-user access. The received SINR and achievable rate expressions for these techniques are provided in the following.
@@ IV-C1 Multi-User Access
In NOMA, spectral resources (i.e., OFDM subcarriers) are fully shared by every user in a cluster and downlink transmissions for different users are multiplexed in the power domain based on superposition coding. In principle, for each cluster, the superposed signal is broadcast to all the involved users and successive interference cancellation (SIC) is applied at the receiver side for multi-user detection. For each user, SIC is performed by detecting and canceling the interference caused by other users with weaker channel conditions while treating the interference originating from those with stronger channel conditions as noise.
@@
Without loss of generality, suppose that the channel gains of the UEs associated with the $u$th cluster are sorted in an ascending order such that:
@@
for $u=1,2,\dots,N_{\mathrm{c}}$. Let $a_{k}^{\mathcal{C}_{u}}$ be the optical power allocation coefficient corresponding to the $k$th UE in the $u$th cluster. The total power constraint implies that $\sum\limits_{k=1}^{K_{u}}\left(a_{k}^{\mathcal{C}_{u}}\right)^{\!2}=1$. To maintain fairness among UEs, a suitable power allocation policy is one that prioritizes the UEs that have weaker channel conditions. In this case, it follows from (46) that $a_{1}^{\mathcal{C}_{u}}\geq\cdots\geq a_{k}^{\mathcal{C}_{u}}\geq\cdots\geq a_{K_{u}}^{\mathcal{C}_{u}}$. The power allocation coefficients are chosen as $a_{k}^{\mathcal{C}_{u}}=\sqrt{\dfrac{K_{u}-k+1}{\sigma}}$, where $\sigma=\dfrac{K_{u}(K_{u}+1)}{2}$ [34]. Upon removing the DC term, the received signal at the $j$th receiver element of the $k$th UE in the $u$th cluster is expressed as (47), given at the top of the page, where the message symbols $X_{\ell}^{\mathcal{C}_{v}}$ are assumed to be drawn from a normalized QAM constellation, and the factor $\zeta=\sqrt{\dfrac{N}{N-2}}$ ensures that the average power of the time domain signal is normalized to unity. Also, the noise terms $V_{kj}$ are independent Gaussian random variables with a zero mean and variance $\sigma_{kj}^{2}$, and the factor $\xi=\dfrac{N-2}{N}$ is the subcarrier utilization ratio. The received signals, for $j=1,2,\dots,7$, are combined together in the form of $Y_{k}^{\mathcal{C}_{u}}=\sum\limits_{j=1}^{7}w_{kj}Y_{kj}^{\mathcal{C}_{u}}$, where $w_{kj}$ is the corresponding combining weight. The receiver fully decodes the SIC term, and then subtracts its remodulated version from the received signal before proceeding to decode the $k$th UE’s message. During the whole decoding process, the receiver treats both the residual multi-user interference (MUI) in the same cluster and the ICI from the rest of the clusters as noise. Assuming perfectly error-free SIC operations, the SINR of the $k$th UE in the $u$th cluster is given by (48) at the top of the page, where $\gamma=\dfrac{\xi^{2}}{R_{\mathrm{PD}}^{2}P_{\mathrm{elec}}}$. Based on MRC, the combining weight for each receiver element is proportional to its received photocurrent to noise ratio [35], resulting in (49) at the top of the page.
When using OFDMA, a fraction $b_{k}^{\mathcal{C}_{u}}$ of the total bandwidth and a fraction $p_{k}^{\mathcal{C}_{u}}$ of the total power are allocated to the $k$th UE in the $u$th cluster subject to $\sum\limits_{k=1}^{K_{u}}b_{k}^{\mathcal{C}_{u}}\!=\!1$ and $\sum\limits_{k=1}^{K_{u}}p_{k}^{\mathcal{C}_{u}}\!=\!1$. Hence, under OFDMA, the intra-cluster MUI is eliminated and the ICI remains as the residual interference. The SINR of the $k$th UE in the $u$th cluster can be written as:
@@ IV-C2 Sum Rate Analysis
and $\Gamma=-\dfrac{\ln\left(5\ \!\mathrm{BER}\right)}{1.5}$ is the SNR gap accounting for the required BER performance.
The achievable rate of the $k$th UE in the $u$th cluster under NOMA and OFDMA is deduced from (52) as:
@@
(a) Borders of SDMA clusters ($N_{\mathrm{c}}=225$)
(b) SINR
(c) Achievable rate
Figure 9: Baseline SDMA clustering and the resulting spatial distributions of the received SINR and data rate on the receiver plane.
V Numerical Results and Discussions
In this section, we evaluate the proposed double-tier AP design for an indoor multi-user access network based on static clustering. The simulation setup consists of an indoor environment as depicted in Fig. 1 with dimensions of $5\times 5\times 3$ m3. The remaining simulation parameters are listed in Table I, unless otherwise specified. Assuming that the receiver is oriented upwards, it requires a half-angle FOV of at least $40^{\circ}$ to be able to receive signals in regions close to the walls. To this end, the half-angle FOV of the ADR is assumed to be $\mathrm{FOV}=50^{\circ}$. The transmit optical power of $P_{\mathrm{t}}=10$ mW per VCSEL is chosen on account of eye safety for each transmitter element based on International electrotechnical commission (IEC) 60825-1 and American national standards institute (ANSI) Z136.1 standards [38, 39, 40]. The full treatment of eye safety is beyond the scope of this paper. In an earlier study [22], we have developed an algorithm for eye safety analysis of a VCSEL array combined with a lens, which computes the maximum transmit power per VCSEL for this purpose.
V-A Clustering Scenarios: Spatial Distribution of Data Rate
In the baseline clustering scenario, every single beam is considered to be carrying independent data for its coverage spot. In this case, clusters are of size one, creating a pure space division multiple access (SDMA) network with the highest possible spatial reuse of the spectral resources. Since the bandwidth is fully reused across the network, the adjacent beams interfere with one another. Fig. LABEL:sub@Fig:6_NoClustering_a illustrates the approximate borders of the beam spots on the receiver plane. Fig. LABEL:sub@Fig:6_NoClustering_b demonstrates the spatial distribution of the received SINR on the same plane. The maximum SINR of $25$ dB is realized at the centers of the beam spots and the SINR value is on the order of $7$ dB at the edge of the network, i.e., in proximity to the walls. To obtain the spatial distribution of the achievable rate, the SINR values are mapped onto their respective data rates based on (53) by assuming that each cluster serves a single [^3] Note that when there is a single UE per cluster, achievable rates for NOMA and OFDMA are the same according to (53). UE. This hypothetical assumption helps to understand the performance bounds in different locations over the coverage area of the AP. According to Fig. LABEL:sub@Fig:6_NoClustering_c, data rates are about $10$ Gb/s to $15$ Gb/s at the centers of the beam spots except in locations near the walls where still a data rate of $8$ Gb/s is provided. For the baseline SDMA, the total number of clusters equals the total number of VCSELs in the AP, i.e., $N_{\mathrm{c}}=225$.